# Sine laws on semigroups with an involutive anti-automorphism: A Levi–Civita approach via left translations

### v4

@@ 5.3. Example 3: $SO(3,\mathbb{R})$ with $\sigma(g)=g^{-1}=g^{\top}$ and column functions
Left action on $V$. Since $(yx)_{i1}=\sum_{j=1}^{3}y_{ij}\,x_{j1}$, we get for $y,x\in SO(3,\mathbb{R})$,
$(L(y)f_{i})(x)=f_{i}(yx)=(yx)_{i1}=\sum_{j=1}^{3}y_{ij}\,f_{j}(x).$
@@
The examples above show concretely how Lemma 3.1 converts the family $R(\sigma(y))$ into left translations $L(y)$ via conjugation by $J$, and how this mechanism can be checked explicitly on natural finite-dimensional spaces of coordinate functions.
@@ 6. Concluding Remarks
The conjugation identity $J\,R(\sigma(y))\,J=L(y)$ provides the key link needed to transfer the anti-automorphic right-translate action into left translations, thereby avoiding the order-reversal obstruction in the family $R(\sigma(\cdot))$. This left-translation viewpoint yields a clean Levi–Civita closure principle in the anti-automorphic setting and leads to analogues of Stetkær’s structural results for the generalized sine law, including the dichotomy $\beta\in\{\pm 1\}$ and the induced $\sigma$-parity rule for $f$. Moreover, under a mild additional centrality hypothesis (e.g. for commutative semigroups), one also recovers the full $\sigma$-transformation rule for $g$. The examples illustrate the conjugation mechanism concretely on standard semigroups and show how the relevant finite-dimensional spaces of coordinate functions behave under left translations.

### v5

@@ 5.3. Example 3: SO(3,R) and column functions
Left action on $V$. Since $(yx)_{i,\,1}=\sum_{j=1}^{3}y_{i,\,j}\,x_{j,\,1}$, we get for $y,x\in SO(3,\mathbb{R})$,
$(L(y)f_{i})(x)=f_{i}(yx)=(yx)_{i,\,1}=\sum_{j=1}^{3}y_{i,\,j}\,f_{j}(x).$
@@
5.4. Example 4: A non-central bridge instance on $S_{3}$
Let $S=S_{3}$ and $\sigma(\pi)=\pi^{-1}$. Define
$g(\pi):=\mathbf{1}_{A_{3}}(\pi),\qquad f(\pi):=\mathbf{1}_{S_{3}\setminus A_{3}}(\pi),$
where $A_{3}$ denotes the alternating subgroup. Then $f$ and $g$ are linearly independent, and for all $x,y\in S_{3}$ one has
$f(x\sigma(y))=f(xy^{-1})=f(x)g(y)+g(x)f(y).$
Indeed, $xy^{-1}$ is odd if and only if $x$ and $y$ have opposite parity. Thus (4.1) holds with $\beta=1$ and $\gamma=0$.
Choose any transposition $d\in S_{3}$. Then $f(d)=1$, and left multiplication by an odd permutation reverses parity, so
$(L(d)f)(x)=f(dx)=g(x)\qquad(\forall x\in S_{3}).$
Hence
$f(d)\neq 0\qquad\text{and}\qquad L(d)f\in\mathrm{span}\{f,g\},$
so the bridge hypothesis (4.6) holds.
On the other hand, $Z(S_{3})=\{e\}$ and $f(e)=0$. Therefore no central element $c$ satisfies $f(c)\neq 0$. This shows that the bridge hypothesis of Theorem 4.6 is strictly weaker than the central-element condition in Corollary 4.7.
The examples above show concretely how Lemma 3.1 converts the family $R(\sigma(y))$ into left translations $L(y)$ via conjugation by $J$, how this mechanism can be checked explicitly on natural finite-dimensional spaces of coordinate functions, and why the intrinsic bridge hypothesis in Theorem 4.6 is the right level of generality for the structural argument.
@@ 6. Concluding Remarks
The conjugation identity $J\,R(\sigma(y))\,J=L(y)$ provides an intrinsic operator interpretation of the order-reversing family $R(\sigma(\cdot))$ that arises from an involutive anti-automorphism. Although an equation of the form $f(x\sigma(y))=\cdots$ can also be rewritten by the substitution $y\mapsto\sigma(y)$, the formula $J\,R(\sigma(y))\,J=L(y)$ shows directly how the anti-automorphic parameter family is converted into a genuine left-translation action. This left-translation viewpoint yields a natural Levi–Civita closure principle in the anti-automorphic setting and leads, under the bridge hypothesis used in Theorem 4.6, to analogues of Stetkær’s structural results for the generalized sine law, including the dichotomy $\beta\in\{\pm 1\}$, the standard $xy$-addition law, and the expected $\sigma$-transformation rules. In the proof, the parity rule $f\circ\sigma=\beta f$ is obtained before the bridge step, whereas the bridge hypothesis is used to control $g\circ\sigma$ and thereby to pass from the $x\sigma(y)$-law to the standard $xy$-addition law. Corollary 4.7 shows that the existence of a central element $c$ with $f(c)\neq 0$ is a convenient sufficient condition for the bridge hypothesis, while Example 5.4 shows that this sufficient condition is not necessary. The examples illustrate the conjugation mechanism concretely on standard semigroups and show how the relevant finite-dimensional spaces of coordinate functions behave under left translations.
Acknowledgment
The author is deeply grateful to the handling editor for the careful handling of the manuscript and to the anonymous referee for reading the paper thoroughly and for providing valuable comments and suggestions, which led to a significant improvement in the presentation.